Standard Quantities Allowed of Labor and Materials

Dulce Company produces chocolate. Each pound of chocolate requires 400 cocoa beans (the unit quantity standard) and 0.15 labor hour (the unit labor standard). During the year, 480,000 pounds of chocolate were produced.

Required:

1. Calculate the number of cocoa beans allowed for the actual output.
beans

2. Calculate the total amount of labor hours allowed for the actual output.
hours
Answer #1

Solution 1:

Number of cocoa beans allowed for the actual output = Actual Output pounds * Cocoa beans per unit

= 480,000 * 400 = 192,000,000 beans

Solution 2:

Labor hours allowed for the actual output = 480,000*0.15 = 72,000 hours